OpenStudy (anonymous):

The diameter of a hydrogen atom is 212pm. Find the length in kilometers of a row of 6.02×1023 hydrogen atoms

OpenStudy (anonymous):

You arrange all of the 6.02x10^23 hydrogen atoms one next to the other in a straight line, so its length will be a sum of the diameters of all atoms. The diameter is given in picometers, where 1 pm = 10^-12 m = 10^-15 km, which means that the diameter of one hydrogen atom in kilometers is 212x10^-15. You multiply that by the number all atoms in the line and you get\[L = 6.02*10^{23}*212*10^{-15}=1276.24*10^{8}=1,27*10^{11} km\] or 127 billion kilometers in simple terms. Quite large a number.